OpenStudy (anonymous):

Divide 3 1/9 by 2 1/9

OpenStudy (anonymous):

28/9 divided 19/9 ->28/9 multiplies 9/19 -> get answer

OpenStudy (anonymous):

Put both into improper fractions. Giving 28/9 and 19/9. Then you get (28/9)/(19/9) Multiply the top by the bottom and the two middle ones together. 252/171 . Goes to 28/19.

OpenStudy (anonymous):

so its 532/81 correct?

OpenStudy (anonymous):

can 532/81 be simplify?

OpenStudy (anonymous):

No. Its 28/19. \[\frac{ 28 }{ 9 }\div \frac{ 19 }{ 9 } = \frac{ 28 }{ 9 } \times \frac{ 9 }{ 19 } = \frac{ 28 }{19 }\]